OpenStudy (anonymous):

What are the zeros of y = 2x^2 + 3x - 1? (Find out what values of x make y = 0.)

OpenStudy (anonymous):

The solutions of this need to be found with the quadratic formula

OpenStudy (anonymous):

Can you show me?

OpenStudy (inkyvoyd):

Do you know the quadratic formula?

OpenStudy (inkyvoyd):

It is \(\huge \frac{-b \pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

Yes, I remember that now.

OpenStudy (inkyvoyd):

Where the quadratic equation is written in the form \(\huge ax^2+bx+c\) Just plug the values in, and see what you get :)

OpenStudy (anonymous):

To make it easier, your equation is in the form \[y=ax^2+bx+c\] So replace a, b, c with numbers.

OpenStudy (anonymous):

put it in quadratic form. y=ax^2 + bx + c

OpenStudy (anonymous):

a=2 b=3 c= -1

OpenStudy (anonymous):

you need to factor it, and set them equal to zero.

OpenStudy (anonymous):

That's the problem. I'm having trouble factoring this equation.

OpenStudy (inkyvoyd):

@hiimchristina , it's not factorable.

OpenStudy (anonymous):

Yeah, I see that..I'm not sure if it is factorable.

OpenStudy (inkyvoyd):

Cam, do what we said, and use the quadratic equation. It's probably the easist way.

OpenStudy (anonymous):

Yeah, it's not... the only factors of 1 are 1 & 1. They can't combine to make 3.

OpenStudy (inkyvoyd):

btw, if the discriminant b^2-4ac is not a perfect square, then it won't come out to rationally factorable.

OpenStudy (anonymous):

OK, I'll do the quadratic equation. :)

OpenStudy (anonymous):

Inside the radical, I get 17. It is not a perfect square.

OpenStudy (anonymous):

OK, I'm sort of stuck now. :( What is the next step?

OpenStudy (anonymous):

\[2x^2 + 3x - 1=0\] \[x=\frac{-3\pm\sqrt{3^2-4\times 2\times -1}}{2\times 2}\] \[=\frac{-3\pm\sqrt{9+8}}{4}=\frac{-3\pm\sqrt{17}}{4}\] there is noting more you can do

OpenStudy (anonymous):

So is the answer "No solution"?

OpenStudy (anonymous):

Or did I misunderstand something?

OpenStudy (anonymous):

But is that what I write as my answer, or is it "No solution"?

OpenStudy (anonymous):

that is the solution

OpenStudy (anonymous):

The question asks what the zeros are. But if there are none, you'd put "None" as the answer, right?

OpenStudy (anonymous):

Either that or there is a misunderstanding...

OpenStudy (anonymous):

I know that, hence the parenthesis in my question. :) I am just trying to clarify that the answer is "None" since the equation isn't even factorable.

OpenStudy (anonymous):

... I'm just going with my answer: "None." I don't want to waste any more time. (No offense.)

OpenStudy (anonymous):

both @Luis_Rivera and i wrote the answer. not sure why you would think there is no answer since we both got the same one

OpenStudy (anonymous):

Read the question. Does your answer answer the question? You're going to have to plug that in, see what happens. :(

OpenStudy (anonymous):

For the value of x.

OpenStudy (anonymous):

:D Well... Maybe I'll just try asking again later. I have 2 more questions to ask today, I don't want them all taking as long as this one.